# Multivariable Behavioral Change Modeling of Epidemics in the Presence of Undetected Infections

### v2

@@ 5 Discussion
In this work we have extended Bayesian behavioral change epidemic models in three important ways to make them more realistic and accurate. We expanded the SIR compartmental structure to incorporate additional data sources on deaths and hospitalizations in an SIHRD models. To the best of our knowledge, this is the first instance of an expanded compartmental structure in a fully stochastic population-averaged model, where both the parameters associated with transition probabilities and all transitions themselves are given distributions and are imputed when not observed. We investigated the incorporation of undetected infections by assuming cases are observed according to a constant detection probability and the true number of infections is estimated using data-augmented MCMC. Finally, we extended the behavioral change framework of Ward et al., (2023) to allow population alarm to be influence by multiple data sources. The proposed model specification also allows us to quantify the relative importance of each data source, so it can be used to investigate potential changes in importance over time.
@@
Our simulation study illustrate several important features of fully Bayesian behavioral change models. First, the incorporation of undetected infections had a minimal impact on the estimation of the effective reproductive number. We believe this is primarily due to the large population size in our simulations, and that this incorporation would be more impactful when modeling epidemics in smaller populations. We did not evaluate this explicitly as our motivation was developing more accurate COVID-19 models in large populations, such as cities like Miami and Montréal. Secondly, we showed that the SIHRD model offers better estimation of $\mathcal{R}_{0}$ and the relative importance parameter in the multivariable alarm, illustrating the increased model complexity is beneficial. Finally when deaths are highly important in influencing behavior through the population alarm, it is critical that deaths are incorporated into the alarm function, particularly for estimating the reproductive number towards the end of the epidemic. However, when deaths are equally as important as cases or minimally important, both the estimation of $\mathcal{R}_{0}$ and overall model fit are comparable between an alarm based on cases only or the more flexible multivariable alarm.
@@
Through analysis of multiple waves of COVID-19 in Miami and Montréal, we illustrate the benefit of our modeling approach in practice. Ignoring behavioral change completely resulted in much poorer model fit than the behavioral change models. We also found that cases were highly important in informing population behavior during the first wave of COVID-19 in both cities, but during the second wave in Miami, deaths were the primary influencing factor of behavior. This finding offers important insight into human behavioral change in response to a multiwave epidemic and may be due to perception going into Wave 2 that the virus was not as deadly as previously thought (Williams et al.,, 2021). The WAIC results from these analyses also indicated that there was minimal downside to using the more complex multivariable alarm function when only cases were important, and a large upside to including deaths when deaths were important. This is meaningful for using these models in practice as we don’t know which data source is influencing behavior prior to model fitting.
In this work we have considered a one-parameter growth function to describe the population alarm, but other functional forms could also be used. Functions used to describe the alarm must satisfy the following properties: for all $x$, $f(x)\in[0,1]$, $f(0)=0$, and for $x\leq y$, $f(x)\leq f(y)$. In words, the alarm can only take values between 0 and 1, must be 0 when there is no disease in the population, and is monotonically increasing as disease incidence increases. In addition to the one-parameter growth function we considered here, Ward et al., (2023) considered two other parametric functions which satisfy this criteria. A two-parameter constant change point model $f(x)=\delta\mathds{1}(x>H)$ with $H\in(0,\text{max}(x))$ $\delta\in[0,1]$, and a three-parameter modified Hill equation (Gesztelyi et al.,, 2012) $f(x)=\frac{\delta}{1+(x_{0}/x)^{\nu}}$ with $\delta\in[0,1]$, $x_{0}\in(0,\text{max}(x))$, and $\nu>0$ which can have a logistic looking shape to represent a smoothed version of the change point function. In considering multivariable extensions of these parametric functions, one must carefully think about identifiability of the overall alarm at any given time point. Both the change point and Hill functions rely on parameters estimating an asymptote ($\delta$) and a mid-point ($H$ or $x_{0}$) indicating the value of the observed data where the alarm is increasing. While this allows these functions to be more flexible, these parameterizations lead to identifiability issues when considering multiple data sources as the same epidemic could occur under various specifications of when the alarm components associated with cases or deaths “went off” (reaches the asymptote) and what level the alarm reaches based on each source (value of the asymptote). For example, an alarm fully attributed to cases may go off once smoothed incidence reaches 100 cases per day, but could also be fully attributed to deaths which reached three deaths at the same time. Or the alarm could be equally attributed to cases and death reaching these thresholds. This is problematic as all three scenarios could lead to the same alarm value as a function of epidemic time.
Alternatively, one may consider non-parametric approaches, such as a multivariate Gaussian process or combination of spline functions. However, care must be taken to ensure the final alarm is constrained between 0 and 1, which can be complicated when using multiple inputs. Additionally, one advantage of the current approach is the interpretability of the $\alpha$ parameter as the relative importance of one data source informing population alarm and its not obvious how to translate this interpretation to a non-parametric approach. Another possible approach would be the use of Bayesian model averaging (Hoeting et al.,, 1999), where each model contains a more complex alarm function based on a single data source and the posterior probabilities of each model would provide the relative importance of the data sources. We plan to pursue both the non-parametric and model averaging approaches as future work to extend this methodology.
@@
Our SIR and SIHRD models have a few limitations that warrant discussion. First, they assume individuals are immediately infectious. While an Exposed compartment can be added to the behavioral change framework as in Ward et al., (2023), this requires strong priors on the latent period and additional data augmentation. Because the SIHRD model already augments several transitions and undetected infections, we omitted the latent period to reduce computational complexity. Second, we do not allow transmission from hospitalized individuals. Although a hospital transmission term could be added to the transmission probability, such effects are typically small, likely unidentifiable from available data, and would complicate interpretation of the alarm function. Third, we assume exponentially distributed infectious and hospitalization periods for the SIHRD model. More flexible distributions could be incorporated using the ‘linear chain trick’ to convolve multiple exponential distributions (Cushing,, 1977), however, this increase in model complexity would likely not have a substantial change on the inference on the reproductive number given that strong priors would be required.

### v3

@@ 5 Discussion
In this work we have extended Bayesian behavioral change epidemic models in three important ways to make them more realistic and accurate. We expanded the SIR compartmental structure to incorporate additional data sources on deaths and hospitalizations in an SIHRD models. To the best of our knowledge, this is the first instance of an expanded compartmental structure in a fully stochastic population-averaged model, where both the parameters associated with transition probabilities and all transitions themselves are given distributions and are imputed when not observed. We investigated the incorporation of undetected infections by assuming cases are observed according to a constant detection probability and the true number of infections is estimated using data-augmented MCMC. Finally, we extended the behavioral change framework of Ward et alward2023bayesian to allow the population alarm to be influenced by multiple data sources. The proposed model specification also allows us to quantify the relative importance of each data source in informing population behavior, so it can be used to investigate potential changes in importance over time.
@@
Our simulation study illustrate several important features of fully Bayesian behavioral change models. We showed that the SIHRD model offers better estimation of $\mathcal{R}_{0}$ and the relative importance parameter in the multivariable alarm, illustrating the increased model complexity is beneficial. We also found that when deaths are highly important in influencing behavior through the population alarm, it is critical that deaths are incorporated into the alarm function, particularly for estimating the reproductive number towards the end of the epidemic. However, when deaths are equally as important as cases or minimally important, both the estimation of $\mathcal{R}_{0}$ and overall model fit are comparable between an alarm based on cases only or the more flexible multivariable alarm.
@@
Through analysis of multiple waves of COVID-19 in Miami and Montréal, we illustrate the benefit of our modeling approach in practice. Ignoring behavioral change completely resulted in much poorer model fit than any of the behavioral change models we considered. We also found that cases were highly important in informing population behavior during the first wave of COVID-19 in both cities, but during the second wave in Miami, deaths were the primary influencing factor of behavior. This finding offers important insight into human behavioral change in response to a multiwave epidemic and may be due to perception going into Wave 2 that the virus was not as deadly as previously thought (williams2021lessons). The WAIC results from these analyses also indicated that there was minimal downside to using the more complex multivariable alarm function when only cases were important, and a large upside to including deaths when deaths were important. This is meaningful for using these models in practice as we don’t know which data source is influencing behavior prior to model fitting.
Our data analysis also illustrated an important limitation of the current modeling framework. In the posterior predictive assessment of the SIHRD model fits for Miami during Wave 2, we found that the model adequately described the observed hospitalization and death trajectories, but the posterior predictive distribution of cases did not match well with the observed data. We hypothesize that this is due to assumption of exponentially distributed infectious and hospitalization periods for the SIHRD model, coupled with the use of strong priors which are needed for identifiability of the undetected infections as described in Section 2.4. More flexible distributions could be incorporated using the ‘linear chain trick’ to convolve multiple exponential distributions (cushing1997integrodifferential), potentially in combination with strong priors informed by previously published literature, e.g. studies of the length-of-stay in the hospital for COVID-19 patients(wen2022time). Our implementation also assumes that the hospitalization rate is constant over the time period modeled, which may not be true as prior research has shown the COVID-19 hospitalization rate to vary over time (couture2022estimating). Our modeling framework would be amenable to varying the hospitalization rate over time, however, strong priors would still be required to incorporate undetected infections. Given that our primary interest was on describing and estimating behavioral change and that our data analysis focused on modeling relatively short time periods (six to ten weeks), we feel that the assumption of constant hospitalization rate over time is reasonable and the violation of this assumption may not be as impactful as it would be for modeling a longer duration.
@@
Our SIR and SIHRD models have a few additional limitations that warrant discussion. First, they assume individuals are immediately infectious. While an Exposed compartment can be added to the behavioral change framework as in Ward et alward2023bayesian, this requires strong priors on the latent period and additional data augmentation. Because the SIHRD model already augments several transitions and undetected infections, we omitted the latent period to reduce computational complexity. Second, we do not allow transmission from hospitalized individuals. Although a hospital transmission term could be added to the transmission probability, such effects are typically small, likely unidentifiable from available data, and would complicate interpretation of the alarm function. Finally, we assume a constant detection rate within the modeled time period, while in reality studies indicate that COVID-19 case ascertainment varied over time (russell2020reconstructing; irons2021estimating). One way to address this is to model multiple waves of relatively short time periods separately, as we did in our data analysis. Incorporating changes in detection within a wave is more challenging when we are simultaneously estimating changes in transmission over time due to behavioral changes, due to the partial identifiability of these models. To overcome this challenge, one could try to incorporate auxiliary data sources such as wastewater (mcmahan2021covid) or seroprevalence surveys (irons2021estimating), and this is an avenue for future work for Bayesian behavioral change models.
Here we have considered a one-parameter growth function to describe the population alarm, but other functional forms could also be used. Functions used to describe the alarm must satisfy the following properties: for all $x$, $f(x)\in[0,1]$, $f(0)=0$, and for $x\leq y$, $f(x)\leq f(y)$. In words, the alarm can only take values between 0 and 1, must be 0 when there is no disease in the population, and is monotonically increasing as disease incidence increases. In addition to the one-parameter growth function we considered here, Ward et alward2023bayesian considered two other parametric functions which satisfy this criteria. A two-parameter constant change point model $f(x)=\delta\mathds{1}(x>H)$ with $H\in(0,\text{max}(x))$ $\delta\in[0,1]$, and a three-parameter modified Hill equation (gesztelyi2012hill) $f(x)=\frac{\delta}{1+(x_{0}/x)^{\nu}}$ with $\delta\in[0,1]$, $x_{0}\in(0,\text{max}(x))$, and $\nu>0$ which can have a logistic looking shape to represent a smoothed version of the change point function. In considering multivariable extensions of these parametric functions, one must carefully think about identifiability of the overall alarm at any given time point. Both the change point and Hill functions rely on parameters estimating an asymptote ($\delta$) and a mid-point ($H$ or $x_{0}$) indicating the value of the observed data where the alarm is increasing. While this allows these functions to be more flexible, these parameterizations lead to identifiability issues when considering multiple data sources as the same epidemic could occur under various specifications of when the alarm components associated with cases or deaths “went off” (reaches the asymptote) and what level the alarm reaches based on each source (value of the asymptote). For example, an alarm fully attributed to cases may go off once smoothed incidence reaches 100 cases per day, but could also be fully attributed to deaths which reached three deaths at the same time. Or the alarm could be equally attributed to cases and death reaching these thresholds. This is problematic as all three scenarios could lead to the same alarm value as a function of epidemic time.
Alternatively, one may consider non-parametric approaches, such as a multivariate Gaussian process or combination of spline functions. However, care must be taken to ensure the final alarm is constrained between 0 and 1, which can be complicated when using multiple inputs. Additionally, one advantage of the current approach is the interpretability of the $\alpha$ parameter as the relative importance of one data source informing population alarm and its not obvious how to translate this interpretation to a non-parametric approach. Another possible approach would be the use of Bayesian model averaging (hoeting1999bayesian), where each model contains a more complex alarm function based on a single data source and the posterior probabilities of each model would provide the relative importance of the data sources. We plan to pursue both the non-parametric and model averaging approaches as future work to extend this methodology.